Suppose you know that a company’s stock currently sells for $67 per share and the required return on the stock is 10.8 percent. You also know that the total return on the stock is evenly divided between capital gains yield and dividend yield. If it’s the company’s policy to always maintain a constant growth rate in its dividends, what is the current dividend per share?  (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)


Dividend Per Share _____________
Answer #2

We know the stock has a required return of 10.8 percent, and the dividend and capital gains yield are equal, so:

Dividend yield = 1/2(.108) = .054 = Capital gains yield

Now we know both the dividend yield and capital gains yield. The dividend is simply the stock price times the dividend yield, so:

D1 = .054($67)

D1 = $3.62

This is the dividend next year. The question asks for the dividend this year. Using the relationship between the dividend this year and the dividend next year:

D1 = D0(1 +g)

$3.62 = D0(1 + .054)

D0 = $3.62 / 1.054

D0 = $3.43


answered by: Nick Lubiani
Answer #1

Required Rate = Dividend Yield + Capital Gain Yield

Dividend Yield = 0.108/2 = 5.40%

Dividend = 0.054(67) = $3.62